OpenStudy (anonymous):

if limit as x approaches 3 is (3x-2)=7, find delta such that (3x-2)-7<.003 whenever 0

OpenStudy (anonymous):

oh, not epsilon delta problem

OpenStudy (anonymous):

yeah i dont understand it :/

OpenStudy (anonymous):

hold on , let me find my Calc text

OpenStudy (anonymous):

thanks

OpenStudy (zarkon):

3x-2-7=3x-9=3(x-3)

OpenStudy (anonymous):

yeh i got that far...but then my book goes into all this stuff i dont understand talking about why the answer (for their example) works

OpenStudy (zarkon):

3(x-3)<.003 x-3<.001

OpenStudy (anonymous):

thats the answer?? i hope so cuz thats what i got....

OpenStudy (zarkon):

Delta=.001

OpenStudy (anonymous):

thank you guys so much

OpenStudy (anonymous):

thank you!

myininaya (myininaya):

you are just using the definition of a limit just plug in the stuff into the definition and woolah

OpenStudy (anonymous):

Basically you know \[|f (x)\to L|<\epsilon \] you wnat \[0<|x-c|<\delta \] so just play with algebra

myininaya (myininaya):

the arrow suppose to be -?

OpenStudy (anonymous):

minus

myininaya (myininaya):

the arrow suppose to be -?

myininaya (myininaya):

:)

myininaya (myininaya):

the arrow suppose to be -?

myininaya (myininaya):

\[\lim_{x \rightarrow c}f(x)=L\]

myininaya (myininaya):

the arrow suppose to be -?

myininaya (myininaya):

what imran is the formal definition for what i just typed

OpenStudy (anonymous):

thanks guys woooo i think i get it now :)

myininaya (myininaya):

the arrow suppose to be -?

myininaya (myininaya):

has*

OpenStudy (anonymous):

Trust me everyone hate these while learning them

OpenStudy (anonymous):

you could say that again....my calc teacher gave us these summer assignments and learning them by myself is the hardest thing

OpenStudy (anonymous):

summer assignment sucks